OpenStudy (anonymous):

Given a cone with a volume of 288 and height 7 in., find the base radius of the cone. Use 3.14 for pi. Round your answer to the tenths place. A. 2.4 b. 2.8 c. 3.5 d. 6.3 i was thinking D

OpenStudy (igreen):

\(V = \dfrac{1}{3} \pi r^2 h\) Plug in what we know: \(288 = \dfrac{1}{3}(3.14)(r^2)(7)\)

OpenStudy (igreen):

Yep, I got 6.3 too.

OpenStudy (igreen):

\(288 = 7.32666667r^2\) \(39.3084622 = r^2\) \(r = \sqrt{39.3084622}\) \(r \approx 6.3\)

OpenStudy (anonymous):

ok thank you. i just wasnt sure bc when i plugged 6.3 in for the radius, i was getting 290. i didnt know if that was good enough to assume that 6.3 was correct. but thank you so much @iGreen